# Normalizing self-supervised learning for provably reliable Change Point Detection

### v1

@@ III-A Spectral Normalization preserves the CPD quality
This problem can be boiled down to a standard two-sample hypothesis testing, which is a focus of this study. We will consider the two most common approaches in this setting: the two-sample tests built on kernel-based statistics and the two-sample likelihood ratio-based tests. Below, it is shown that spectral normalization ensures the transition into the space of representations:
@@
    in the case of kernel-based tests — does not change the type II error rate of convergence to zero (for no assumptions about the data distribution at all);
@@
    in the case of likelihood ratio-based tests — does not decrease the power of such tests (for reasonable assumptions about the data distribution).
@@ III-A1 Bi-Lipschitz neural networks
$L_{1}\lVert\mathbf{X}-\mathbf{X}^{\prime}\rVert_{\mathcal{X}}\leq\lVert g_{l}(\mathbf{X})-g_{l}(\mathbf{X}^{\prime})\rVert_{\mathcal{H}}\leq L_{2}\lVert\mathbf{X}-\mathbf{X}^{\prime}\rVert_{\mathcal{X}},$ (5)
@@ III-A3 Kernel-based test
Consider two sequences of observations $\mathbf{X}=[\mathbf{x}_{1},\dots,\mathbf{x}_{n}]$ and $\hat{\mathbf{X}}=[\hat{\mathbf{x}}_{1},\dots,\hat{\mathbf{x}}_{n}],\mathbf{x}_{i},\hat{\mathbf{x}}_{i}\in\mathbb{S}^{D-1}$. Denote by $\mathbf{Y},\,\hat{\mathbf{Y}}$ their images under the distance-preserving mapping $G$. Consider a bounded kernel $0\leq k(\mathbf{x},\mathbf{x}^{\prime})\leq K$ that is preserved under $G$ and the corresponding sample $\mathrm{MMD}^{k}_{b}$ statistic. The $\mathrm{MMD}^{k}_{b}$-based two-sample test of level $\alpha$ has the same type II error rate $O(m^{-\frac{1}{2}})$ of convergence to zero in the space of embeddings $\mathcal{Y}$ as in the space of raw observations $\mathcal{X}$.
@@ IV-E The dynamics of the representations
Figure 6: The dynamics of the representations. Blue color corresponds to vanilla TS2Vec, and orange — to its SN version. Sliding window sizes: 60 for HASC, 75 for Yahoo!A4Benchmark, 100 for USC-HAD. Each value at the position $i$ corresponds to the similarity between te representations of $\mathbf{X}[i:i+w]$ and $\hat{\mathbf{X}}[i:i+w]$.
@@ A-D Proof of Theorem 1
$\mathrm{MMD}_{b}(Z,\Xi)<\sqrt{\frac{2K}{m}}\bigl{(}1+\sqrt{2\log\alpha^{-1}}\bigr{)},$
@@
$\mathbb{P}\bigl{\{}|\mathrm{MMD}_{b}(Z,\Xi)$ $-\mathrm{MMD}_{k}(P,Q)|>4\left(\frac{K}{m}\right)^{\frac{1}{2}}+\varepsilon\bigr{\}}\leq$ (23)
$\leq$ $2\exp{\left(\frac{-\varepsilon^{2}m}{4K}\right)}.$
The equation (23) shows that the type II error converges to zero at rate $O(m^{-1/2})$; note that the change of constant $K$ to another constant $\tilde{C}K$ does not affect the convergence rate. Therefore, for the $\mathrm{MMD}_{b}$-based hypothesis test of level $\alpha$, the type II error convergence rate in the space of embeddings is the same as in the space of observations.

### v2

@@ III-A Spectral Normalization preserves the CPD quality
We will consider the two most common approaches in this setting: the two-sample tests built on kernel-based statistics and the likelihood ratio-based tests. Below, it is shown that spectral normalization ensures the transition into the space of representations:
@@
    in the case of two-sample kernel-based tests — does not change the type II error rate of convergence to zero (for no assumptions about the data distribution at all);
@@
    in the case of likelihood ratio-based CPD tests — does not decrease the power of such tests (for reasonable assumptions about the data distribution).
@@ III-A1 Bi-Lipschitz neural networks
$L_{1}\lVert\mathbf{X}-\mathbf{X}^{\prime}\rVert_{\mathcal{X}}\leq\lVert h(\mathbf{X})-h(\mathbf{X}^{\prime})\rVert_{\mathcal{H}}\leq L_{2}\lVert\mathbf{X}-\mathbf{X}^{\prime}\rVert_{\mathcal{X}},$ (5)
@@ III-A3 Kernel-based test
Consider two sequences of observations $\mathbf{X}=[\mathbf{x}_{1},\dots,\mathbf{x}_{n}]$ and $\hat{\mathbf{X}}=[\hat{\mathbf{x}}_{1},\dots,\hat{\mathbf{x}}_{n}],\mathbf{x}_{i},\hat{\mathbf{x}}_{i}\in\mathbb{S}^{D-1}$. Denote by $\mathbf{Y},\,\hat{\mathbf{Y}}$ their images under the distance-preserving mapping $G$. Consider a bounded kernel $0\leq k(\mathbf{x},\mathbf{x}^{\prime})\leq K$ that is preserved under $G$ and the corresponding sample $\mathrm{MMD}^{k}_{b}$ statistic. The $\mathrm{MMD}^{k}_{b}$-based two-sample test of level $\alpha$ has the same type II error rate $O(n^{-\frac{1}{2}})$ of convergence to zero in the space of embeddings $\mathcal{Y}$ as in the space of raw observations $\mathcal{X}$.
@@ IV-E The dynamics of the representations
Figure 6: The dynamics of the representations. Blue color corresponds to vanilla TS2Vec, and orange — to its SN version. Sliding window sizes: 60 for HASC, 75 for Yahoo!A4Benchmark, 100 for USC-HAD. Each value at the position $i$ corresponds to the similarity between the representations of $\mathbf{X}[i:i+w]$ and $\hat{\mathbf{X}}[i:i+w]$.
@@ A-D Proof of Theorem 1
$\mathrm{MMD}_{b}(Z,\Xi)<\sqrt{\frac{2K}{n}}\bigl{(}1+\sqrt{2\log\alpha^{-1}}\bigr{)},$
@@
$\mathbb{P}\bigl{\{}|\mathrm{MMD}_{b}(Z,\Xi)$ $-\mathrm{MMD}_{k}(P,Q)|>4\left(\frac{K}{n}\right)^{\frac{1}{2}}+\varepsilon\bigr{\}}\leq$ (23)
$\leq$ $2\exp{\left(\frac{-\varepsilon^{2}n}{4K}\right)}.$
The equation (23) shows that the type II error converges to zero at rate $O(n^{-1/2})$; note that the change of constant $K$ to another constant $\tilde{C}K$ does not affect the convergence rate. Therefore, for the $\mathrm{MMD}_{b}$-based hypothesis test of level $\alpha$, the type II error convergence rate in the space of embeddings is the same as in the space of observations.